# A Survey of World Models for Autonomous Driving

### v3

@@ 6.1 Evaluation Platforms and Benchmarks
     OccWorld [85]      3D-Occ      None      25.78      15.14      10.51      17.14      34.63      25.07      20.18      26.63
     RenderWorld [91]      3D-Occ      None      28.69      18.89      14.83      20.80      37.74      28.41      24.08      30.08
     OccLLaMA-O [89]      3D-Occ      None      25.05      19.49      15.26      19.93      34.56      28.53      24.41      29.17
     DOME-O [95]      3D-Occ      None      35.11      25.89      20.29      27.10      43.99      35.36      29.74      36.36
     DFIT-OccWorld-O [86]      3D-Occ      None      31.68      21.29      15.18      22.71      40.28      31.24      25.29      32.27
[…]
     TPVFormer [224]+SelfOcc [211]+OccWorld-S [85]      Camera      None      0.28      0.26      0.24      0.26      5.05      5.01      4.95      5.00
     OccWorld-F [89]      Camera      None      8.03      6.91      3.54      6.16      23.62      18.13      15.22      18.99
     OccLLaMA-F [89]      Camera      None      10.34      8.66      6.98      8.66      25.81      23.19      19.97      22.99
     RenderWorld [91]      Camera      None      2.83      2.55      2.37      2.58      14.61      13.61      12.98      13.73
     DOME-F [95]      Camera      None      24.12      17.41      13.24      18.25      35.18      27.90      23.435      28.84
     DFIT-OccWorld [86]      Camera      3D-Occ      13.38      10.16      7.96      10.50      19.18      16.85      15.02      17.02
TABLE IX: Motion planning on the nuScenes [36] dataset6.5). Aux.Sup. denotes auxiliary supervision apart from the ego trajectory.
@@
      NMP [133]       LiDAR       Box & Motion       0.53       1.25       2.67       1.48       0.04       0.12       0.87       0.34
      FF [134]       LiDAR       Freespace       0.55       1.20       2.54       1.43       0.06       0.17       1.07       0.43
      EO [117]       LiDAR       Freespace       0.67       1.36       2.78       1.60       0.04       0.09       0.88       0.33
[…]
      UniAD [225]       Camera        Map & Box & Motion & Tracklets & Occ       0.48       0.96       1.65       1.03       0.05       0.17       0.71       0.31
      UniAD+DriveWorld [21]       Camera        Map & Box & Motion & Tracklets & Occ       0.34       0.67       1.07       0.69       0.04       0.12       0.41       0.19
      DriveDreamer [10]       Camera       Map & Box & Motion       -       -       -       0.29       -       -       -       0.15
      GenAD [22]       Camera       Map & Box & Motion       0.36       0.83       1.55       0.91       0.06       0.23       1.00       0.43
      OccWorld-T [85]       Camera       Semantic LiDAR       0.54       1.36       2.66       1.52       0.12       0.40       1.59       0.70
      OccWorld-S [85]       Camera       None       0.67       1.69       3.13       1.83       0.19       1.28       4.59       2.02
      Drive-OccWorld [88]       Camera       None       0.32       0.75       1.49       0.85       0.05       0.17       0.64       0.29
      ViDAR [20]       Camera       None       -       -       -       0.91       -       -       -       0.23
      OccWorld-F [89]       Camera       Occ       0.45       1.33       2.25       1.34       0.08       0.42       1.71       0.73
      OccLLaMA-F [89]       Camera       Occ       0.38       1.07       2.15       1.20       0.06       0.39       1.65       0.70
      RenderWorld [91]       Camera       Occ       0.48       1.30       2.67       1.48       0.14       0.55       2.23       0.97
      DFIT-OccWorld-V [86]       Camera       Occ       0.42       1.14       2.19       1.25       0.09       0.19       1.37       0.55
[…]
      RenderWorld [91]       3D-Occ       None       0.35       0.91       1.84       1.03       0.05       0.40       1.39       0.61
      OccLLaMA-O [89]       3D-Occ       None       0.37       1.02       2.03       1.14       0.04       0.24       1.20       0.49
      DFIT-OccWorld-O [86]       3D-Occ       None       0.38       0.96       1.73       1.02       0.07       0.39       0.90       0.45
6.3 Point Cloud Forecasting
Results. Table VII shows that both DFIT‐OccWorld variants substantially outperform the ViDAR baseline at every forecast horizon. DFIT-OccWorld-O trims mean CD to 0.70m2 and stays ahead of ViDAR across all horizons; the V-variant follows at 0.76m2, and the gap widens as lead time grows. The field is shifting from range-image CNN/LSTM to token-based diffusion, leveraging image priors and physics constraints to sustain long-horizon precision.
6.4 4D Occupancy Forecasting
@@ 6.3 Point Cloud Forecasting
Results. Table VIII summarizes the 4D occupancy forecasting performance on Occ3D-nuScenes [83], where predictions for 1s, 2s, and 3s into the future are assessed via mIoU and IoU. Notably, DOME-O attains state-of-the-art results (27.10% mIoU and 36.36% IoU), surpassing baseline methods such as OccWorld and RenderWorld by substantial margins. Even the purely camera-based DOME-F variant remains highly competitive, reflecting the model’s robustness in scenarios without direct 3D occupancy supervision. Future advances target tokenised or diffusion backbones that accept language or trajectory prompts and camera-only pipelines that cut LiDAR cost while keeping real-time viability.
6.5 Motion planning
@@
Results. Table IX presents a quantitative comparison of motion planning methods on the nuScenes [36] dataset. End-to-end planners (e.g., UniAD [225]) achieve the lowest average L2 error and collision rate when trained with rich map, box and motion labels. Occupancy-centric methods (e.g., OccWorld [85], RenderWorld [91], DFIT-OccWorld [86]) retain competitive accuracy while discarding most auxiliary supervision, especially in camera-only setups. RenderWorld further achieves a $\sim$34reduction in collision rate over a 3-second horizon compared to OccWorld, indicating safer long-horizon forecasts. Overall, the table highlights how both occupancy-centric and end-to-end solutions excel at generating precise, collision-free paths, with top-performing models striking an effective balance between minimal supervision and accurate motion forecasting.
7 Future Research Directions
@@
7.1 Self-Supervised World Models
@@
7.2 Multi-Modal World Models
@@
7.3 Advanced Simulation
${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ autono- ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ videos [${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ deployment.
@@ 7.3 Advanced Simulation
7.4 Efficient World Models
${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ slashing computation. Multi-scale grids will stream context before zooming into high-detail regions around dynamic agents, conserving ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ reach.
8 Conclusion

### v4

@@ 6.3. Point Cloud Forecasting
OccWorld (Zheng et al., 2025a) 3D-Occ None 25.78 15.14 10.51 17.14 34.63 25.07 20.18 26.63
RenderWorld (Yan et al., 2025a) 3D-Occ None 28.69 18.89 14.83 20.80 37.74 28.41 24.08 30.08
OccLLaMA-O (Wei et al., 2024) 3D-Occ None 25.05 19.49 15.26 19.93 34.56 28.53 24.41 29.17
Occ-LLM (Xu et al., 2025a) 3D-Occ None 24.02 21.65 17.29 20.99 36.65 32.14 28.77 32.52
DFIT-OccWorld-O (Zhang et al., 2024c) 3D-Occ None 31.68 21.29 15.18 22.71 40.28 31.24 25.29 32.27
DOME-O (Gu et al., 2024) 3D-Occ None 35.11 25.89 20.29 27.10 43.99 35.36 29.74 36.36
AFMWM(Liu et al., 2025) 3D-Occ None 36.42 27.39 21.66 28.49 43.68 36.89 31.98 37.52
UniScene (Li et al., 2024a) 3D-Occ None 35.37 29.59 25.08 31.76 38.34 32.70 29.09 34.84
I2-World-O (Liao et al., 2025) 3D-Occ None 47.62 38.58 32.98 39.73 54.29 49.43 45.69 49.80
T3Former-O (Xu et al., 2025b) 3D-Occ None 46.32 33.23 28.73 36.09 77.00 75.89 76.32 76.40
[1pt/1pt] RenderWorld (Yan et al., 2025a) Camera Occ 2.83 2.55 2.37 2.58 14.61 13.61 12.98 13.73
TPVFormer (Huang et al., 2023b)+Lidar+OccWorld-T (Zheng et al., 2025a) Camera Semantic LiDAR 4.68 3.36 2.63 3.56 9.32 8.23 7.47 8.34
TPVFormer (Huang et al., 2023b)+SelfOcc (Huang et al., 2024b)+OccWorld-S (Zheng et al., 2025a) Camera None 0.28 0.26 0.24 0.26 5.05 5.01 4.95 5.00
OccWorld-F (Wei et al., 2024) Camera Occ 8.03 6.91 3.54 6.16 23.62 18.13 15.22 18.99
OccLLaMA-F (Wei et al., 2024) Camera Occ 10.34 8.66 6.98 8.66 25.81 23.19 19.97 22.99
GWM (Feng et al., 2025) Camera Occ 11.63 10.07 8.17 10.12 26.22 24.97 22.13 24.60
DOME-F (Gu et al., 2024) Camera None 24.12 17.41 13.24 18.25 35.18 27.90 23.435 28.84
DFIT-OccWorld (Zhang et al., 2024c) Camera 3D-Occ 13.38 10.16 7.96 10.50 19.18 16.85 15.02 17.02
I2-World-STC (Liao et al., 2025) Camera None 21.67 18.78 16.47 18.97 30.55 28.76 26.99 28.77
T3Former-F (Xu et al., 2025b) Camera Occ 24.87 18.30 15.63 19.60 38.98 37.45 31.89 36.11
Table 9. Motion planning on the nuScenes (Caesar et al., 2020) dataset6.5). Aux.Sup. denotes auxiliary supervision apart from the ego trajectory. We source the results from their respective papers.
@@
NMP (Zeng et al., 2019) LiDAR Box & Motion 0.53 1.25 2.67 1.48 0.04 0.12 0.87 0.34
FF (Hu et al., 2021a) LiDAR Freespace 0.55 1.20 2.54 1.43 0.06 0.17 1.07 0.43
EO (Khurana et al., 2022) LiDAR Freespace 0.67 1.36 2.78 1.60 0.04 0.09 0.88 0.33
[1pt/1pt] ST-P3 (Hu et al., 2022a) Camera Map & Box & Depth 1.33 2.11 2.90 2.11 0.23 0.62 1.27 0.71
UniAD (Hu et al., 2023b) Camera Map & Box & Motion & Tracklets & Occ 0.48 0.96 1.65 1.03 0.05 0.17 0.71 0.31
UniAD+DriveWorld (Min et al., 2024b) Camera Map & Box & Motion & Tracklets & Occ 0.34 0.67 1.07 0.69 0.04 0.12 0.41 0.19
DriveDreamer (Wang et al., 2024c) Camera Map & Box & Motion - - - 0.29 - - - 0.15
GenAD (Zheng et al., 2025b) Camera Map & Box & Motion 0.36 0.83 1.55 0.91 0.06 0.23 1.00 0.43
OccWorld-T (Zheng et al., 2025a) Camera Semantic LiDAR 0.54 1.36 2.66 1.52 0.12 0.40 1.59 0.70
OccWorld-S (Zheng et al., 2025a) Camera None 0.67 1.69 3.13 1.83 0.19 1.28 4.59 2.02
Drive-OccWorld (Yang et al., 2025c) Camera None 0.32 0.75 1.49 0.85 0.05 0.17 0.64 0.29
ViDAR (Yang et al., 2024a) Camera None - - - 0.91 - - - 0.23
OccWorld-F (Wei et al., 2024) Camera Occ 0.45 1.33 2.25 1.34 0.08 0.42 1.71 0.73
OccLLaMA-F (Wei et al., 2024) Camera Occ 0.38 1.07 2.15 1.20 0.06 0.39 1.65 0.70
RenderWorld (Yan et al., 2025a) Camera Occ 0.48 1.30 2.67 1.48 0.14 0.55 2.23 0.97
DFIT-OccWorld-V (Zhang et al., 2024c) Camera Occ 0.42 1.14 2.19 1.25 0.09 0.19 1.37 0.55
GWM (Feng et al., 2025) Camera Occ 0.24 1.01 2.05 1.13 0.07 0.26 1.45 0.59
UncAD (Yang et al., 2025d) Camera None 0.33 0.59 0.94 0.62 0.10 0.14 0.28 0.17
FSDrive (Zeng et al., 2025) Camera Ego status 0.14 0.25 0.46 0.28 0.03 0.06 0.21 0.10
[1pt/1pt] OccWorld (Zheng et al., 2025a) 3D-Occ None 0.43 1.08 1.99 1.17 0.07 0.38 1.35 0.60
RenderWorld (Yan et al., 2025a) 3D-Occ None 0.35 0.91 1.84 1.03 0.05 0.40 1.39 0.61
OccLLaMA-O (Wei et al., 2024) 3D-Occ None 0.37 1.02 2.03 1.14 0.04 0.24 1.20 0.49
DFIT-OccWorld-O (Zhang et al., 2024c) 3D-Occ None 0.38 0.96 1.73 1.02 0.07 0.39 0.90 0.45
T3Former-O (Xu et al., 2025b) 3D-Occ None 0.32 0.91 1.76 1.00 0.08 0.32 0.51 0.30
6.4. 4D Occupancy Forecasting
@@
Results. Table 8 summarizes the 4D occupancy forecasting performance on Occ3D-nuScenes (Tian et al., 2024), where predictions for 1s, 2s, and 3s into the future are assessed via mIoU and IoU. Notably, I2-World-O (Liao et al., 2025) and T3Former-O (Xu et al., 2025b) attain state-of-the-art results, surpassing baseline methods. Even the purely camera-based T3Former-F variant remains highly competitive, reflecting the model’s robustness in scenarios without direct 3D occupancy supervision. Future advances target tokenised or diffusion backbones that accept language or trajectory prompts and camera-only pipelines that cut LiDAR cost while keeping real-time viability.
6.5. Motion planning
@@
[…]
7. Future Research Directions
@@
7.1. Self-Supervised World Models
@@
7.2. Multi-Modal World Models
@@
7.3. Advanced Simulation
Physics Engine. Physics-aware generation steers autonomous-driving simulators. DrivePhysica aligns ego/world frames, injects 3D flows, and applies box-guided occlusion, yielding multi-camera videos (Yang et al., 2024c). Future platforms fuse diffusion world models with differentiable rigid, soft, fluid, and thermal solvers in one loop. Diffusion priors sketch 4D occupancy while the solver enforces Newtonian laws, ensuring validity and gradients that reveal traction loss, drag, or sensor occlusion before deployment.
@@ 7.3. Advanced Simulation
7.4. Efficient World Models
Looking forward, efficient world models will rely on lean, latency-aware backbones that fuse sensing, prediction, and control in a single sweep. One shared feature reservoir – drawing jointly on vision, depth, and map cues – will serve every downstream head, eliminating costly re-projection. Lightweight language adapters will weave route guidance and traffic regulations into this common space. Scene motion will be roughed out as coarse flow; only ambiguous regions will be up-sampled into dense volumes, slashing computation. Multi-scale grids will stream context before zooming into high-detail regions around dynamic agents, conserving memory. Capacity-elastic layers will monitor feature entropy and activate deeper blocks only when congestion rises, sustaining real-time performance on modest vehicle GPUs. Coupled with edge quantisation and federated updates, these systems will learn continually from fleet data while respecting power, bandwidth, and privacy limits – bringing scalable, safety-critical autonomy within tangible reach.
8. Conclusion